E/x=C+3/x to find c, the number of cartons that were packed.
a. C=E-3/x
b. C=E/x-3
c. C=E/3-x
d. C=3-E/x

Answer by Boreal(15235)

E/x=C+(3/x)
(E/x)-(3/x)=C
(E-3)/x=C
A
The parentheses are critical
B has the wrong denominator
3-E is wrong
It can be (E/x)-(3/x) or what was written above.